OpenStudy (anonymous):

Determine the zeros of the function 3x2 - 8 = 10x.

OpenStudy (anonymous):

options are: A. {-2, 4} B. {2, 4} C. the set of negative two-thirds and 4 D. the set of two-thirds and 4

Directrix (directrix):

The zeros of 3x^2 - 8 = 10x are the values of x that make the polynomial equal to zero. 3x^2 - 8 = 10x 3 x^2 - 10x - 8 = 0 Factor the polynomial.

Directrix (directrix):

3 x^2 - 10x - 8 = 0 (x - ?) (3x + ?) = 0 --> @AlDonnelly

OpenStudy (anonymous):

thank you so much, this helped a lot.